# Language Guided Skill Discovery

### v1

@@ Abstract
Skill discovery methods enable agents to learn diverse emergent behaviors without explicit rewards. To make learned skills useful for unknown downstream tasks, obtaining a semantically diverse repertoire of skills is essential. While some approaches introduce a discriminator to distinguish skills and others aim to increase state coverage, no existing work directly addresses the “semantic diversity” of skills. We hypothesize that leveraging the semantic knowledge of large language models (LLMs) can lead us to improve semantic diversity of resulting behaviors. In this sense, we introduce Language Guided Skill Discovery (LGSD), a skill discovery framework that aims to directly maximize the semantic diversity between skills. LGSD takes user prompts as input and outputs a set of semantically distinctive skills. The prompts serve as a means to constrain the search space into a semantically desired subspace, and the generated LLM outputs guide the agent to visit semantically diverse states within the subspace. We demonstrate that LGSD enables legged robots to visit different user-intended areas on a plane by simply changing the prompt. Furthermore, we show that language guidance aids in discovering more diverse skills compared to five existing skill discovery methods in robot-arm manipulation environments. Lastly, LGSD provides a simple way of utilizing learned skills via natural language.
@@ 1 Introduction
One of the key capabilities of intelligent agents is to autonomously learn useful skills applicable to downstream tasks without task-specific objectives. Consider how human infants acquire manipulation skills through simple play with toys, which can later be applied to a broader set of tasks, such as using a spoon or holding a bottle. Our research aims to emulate this process of skill learning. Prior works in unsupervised skill discovery suggest that maximizing diversity of behaviors can be a way to develop such skills [Gregor et al., 2016, Eysenbach et al., 2018, Kwon, 2020]. These approaches operate under the assumption that acquiring a wide range of diverse behaviors may lead to the development of useful skills for downstream tasks. Building on this premise, studies like Sharma et al. [2019], Hansen et al. [2019], Liu and Abbeel [2021a], Laskin et al. [2022] have associated behaviors with randomly sampled vectors by maximizing Mutual Information between them to acquire diverse skills. Additionally, there are other approaches aimed at increasing coverage of the state space [Liu and Abbeel, 2021b, Yarats et al., 2021, Park et al., 2021], or exploration-based strategies that leverage errors in predictive models to enhance learning [Burda et al., 2018, Pathak et al., 2019, Park et al., 2023a].
@@
However, we contend that these measures are proxies for what we term “semantic diversity”, and existing proxies objects do not necessarily reflect the semantic diversity of a repertoire of skills. For instance, in high degrees of freedom (DOF) systems, such as robot-arm manipulation tasks, unstructured movements of each DOF might exhibit high diversity in terms of state coverage or distinctiveness from a neural discriminator, but nonetheless lack meaningful semantic diversity. To more accurately measure semantic diversity, we propose the use of Large Language Models (LLMs) as they are well-suited for understanding and assessing the semantic significance of different behaviors in a way that better aligns with human judgment.
Figure 1: We proposed LGSD  which can discover a semantically distinctive set of skills. We display four sample skills acquired from a single training run. Red dots indicate the path of the end-effector, while blue dots represent the trajectory of the object moved. Two skills moved the green cube to the right, while others moved the red cube to the top.
@@
In this work, we present Language Guided Skill Discovery (LGSD), a skill discovery method that utilizes the guidance of LLMs to learn semantically diverse skills. LGSD aims to fulfill three desiderata for skill discovery. (i) Firstly, we want to discover skills that have diverse semantic meanings. We use LLMs to generate descriptions for each of agent states. Based on these descriptions, we then measure the semantic difference between states using language-distance, and train skills to maximize this distance. Since LLMs are equipped with a semantic understanding of the world, this provides a way to maximizing semantic differences between skills. Fig. 1 illustrates skills learned from LGSD, showcasing the semantic diversity of resultant behaviors. (ii) Secondly, we seek to restrict the search space for skills into a desired semantic subspace. Intuitively, restricting the skill space is akin to enforcing an infant to grab toys only with their hands, and not with their mouth or feet. We implement this desideratum by utilizing language prompts. Humans provide prompts to LLMs, and the generated descriptions are then constrained by these prompts. This will allow our method to ignore differences outside the intended semantic space. (iii) Lastly, we aim for the learned skills to be easy to use. When the skill space is continuous, merely selecting the best skills for target tasks becomes non-trivial. LGSD supports a zero-shot language instruction-following capability by training separate network $\psi$ for skill inference. Humans can provide a natural language description of the desired state, and then LGSD can infer which skill should be used to reach that goal state.
@@ 2.1 Unsupervised skill discovery
Common approaches to associate skills with corresponding behaviors are based on maximizing the mutual information (MI) between states $S$ and skills $Z$, or $\mathcal{I}(S;Z)$ [Gregor et al., 2016, Eysenbach et al., 2018, Sharma et al., 2019, Hansen et al., 2019, Liu and Abbeel, 2021a, Laskin et al., 2022]. MI can be viewed as a dependency measure that quantifies the association between two random variables. To maximize this quantity, a popular early formulation utilizes the decomposition $\mathcal{I}(S;Z)=\mathcal{H}(Z)-\mathcal{H}(Z|S)$, where $\mathcal{H}[\cdot]$ refers to Shannon entropy. Since directly computing $\mathcal{H}(Z|S)$ is intractable, Eysenbach et al. [2018] introduce a variational posterior $q_{\phi}(z|s)$ to optimize the lower bound. Similarly, CIC [Laskin et al., 2022] maximizes $\mathcal{I}((s,s^{\prime});Z)$, promoting distinct state transitions across different skills. However, a significant limitation of using MI as the objective arises from the use of Kullback–Leibler (KL) divergence to define MI:
@@
Maximization of MI using KL divergence often results in skills with less distinctive behaviors. This is because KL divergence is fully maximized when two densities have no overlap, posing a problem as there is no further incentive to distinguish the densities beyond this point. This is problematic because having no overlap between two densities does not necessarily mean two skills are noticeably distinctive. In practice, this becomes a severe shortcoming of DIAYN [Eysenbach et al., 2018], where its objective can be fully optimized as long as the discriminator $q_{\phi}$ can distinguish them perfectly. Neural networks can easily distinguish minor numerical differences, so the skills learned with this objective often have only these minute differences.
@@
To overcome these shortcomings, recent skill discovery works [He et al., 2022, Park et al., 2021, 2023a, 2023b] propose using the Wasserstein Dependency Measure (WDM,  Ozair et al. [2019]):
@@
An essential characteristic of $\mathcal{I_{W}}$ is that it must be defined within a metric space. Therefore, the choice of metric to measure the difference between skills comes to the gist of the algorithm since it ultimately governs the resulting behavior of each skills. LSD [Park et al., 2021] employs the Euclidean distance in state space as its metric. Maximizing $\mathcal{I_{W}}$ under this metric encourages agents to visit states that are are far apart in terms of Euclidean distances. Conversely, CSD [Park et al., 2023a] points out the problem of using Euclidean distance as the metric. They argue that maximizing Euclidean distance tends to focus on "easy" changes in the state space, such as moving the robot arm itself rather than the target object. Thus, CSD proposed to maximize "controllability-aware distance" instead. This metric measures the negative log likelihood of a transition dynamics $-\log{p(s^{\prime}|s)}$. Using this metric means the algorithm favors visiting rare transitions. LGSD adopts a similar approach but differentiates itself by employing a language-based distance metric within the WDM framework, to maximize the semantic difference.
@@ 2.2 Using language as guidance for behavior learning
Recent work has studied using semantic knowledge in large pre-trained models for control tasks in several ways. A popular approach is to use the common-sense reasoning capabilities of LLMs to produce high-level plans over known, existing skills through the interface of natural language [Sharma et al., 2021, Ahn et al., 2022, Singh et al., 2023, Huang et al., 2022, Ha et al., 2023] or generated programs [Liang et al., 2022, Zeng et al., 2022, Vemprala et al., 2024]. In addition, LLMs have also been shown to be effective in directly producing low-level actions that can be executed by an agent [Wang et al., 2023, Kwon et al., 2023, Driess et al., 2023, Brohan et al., 2023]. Our work does not use LLMs to generate actions at any level of abstraction but to provide a meaningful distance measure to guide learning towards discovering semantically meaningful skills.
Several works have used LLMs to guide agents to learn semantically meaningful behavior by using them to design reward functions for either a given desired task [Yu et al., 2023, Ma et al., 2023]. This showcases LLMs’ capacity to understand specific physical embodiments assorting it with semantic knowledge to reason about given tasks. We note that these works aim to solve a single specific task, whereas our method aims to learn a repertoire of diverse behaviors without explicit supervision.
The work most related to ours is Du et al. [2023], which employs an LLM to guide exploration. Their approach uses the LLM to suggest plausible set of goals for the agent, maximizing the cosine similarity between the language embeddings of the goals and the current state. While they focus on using an LLM to reason with a low-dimensional, discrete set of high-level actions in complex environments like Crafter and Housekeep, our work focus on discovering meaningful behaviors in complex, continuous control spaces where actions (e.g. joint targets) are not as readily interpretable. Furthermore, their reward function, based on language similarity, lacks theoretical support as it is not a valid metric. In contrast, our method effectively handles the language-distance measure and provides a theoretical framework for its use. We elaborate on this in Section 4.3.
@@ 4.1 Algorithm
Our goal is for an agent to learn a corpus of skills that visits semantically different states. Discovering semantically distinct states requires 1) a way of measuring the semantic difference between states, and 2) a means of maximizing this difference measure. To evaluate the semantic differences between different behaviors, we leverage the power of LLMs. We will consider two states as substantially distinct if the language descriptions produced by LLMs are semantically different. As illustrated in Fig. 2, an LLM is queried to produce a natural language description for each state. This description links the state to its semantic meaning. We then measure the difference between these language descriptions. To measure the difference between two language descriptions, we leverage a pre-trained natural language embedding model, Sentence-Transformer [Reimers and Gurevych, 2019], to map the language descriptions into fixed-dimension real-valued vectors. This allows us to measure the difference between these vectors using cosine similarity. We refer to this difference measure as language-distance and denote it as $d_{\text{lang}}$.
@@
Spreading out in the latent space can be achieved by aligning the vector $\phi(s^{\prime})-\phi(s)$ with a latent skill vector $z$ sampled from a Gaussian distribution $z\sim\mathcal{N}(0,I)$. Because the vector $z$ is sampled isotropically, aligning $\phi(s^{\prime})-\phi(s)$ with $z$ fosters a diversity of behaviors depending on the sampled $z$. We use the reward $r(s,z,s^{\prime})=(\phi(s^{\prime})-\phi(s))^{T}z$ and maximize the cumulative reward to achieve this objective. Here, Park et al. [2023b] showed that maximizing the cumulative reward $r(s,z,s^{\prime})$ with a 1-Lipschitz continuous $\phi$ is equivalent to maximizing the WDM between the terminal state $S_{T}$ and $Z$:
$\mathop{\rm argmax}_{\pi,\phi}\ {\mathcal{I_{W}}(S_{T};Z)}$ $\approx\mathop{\rm argmax}_{\pi,\phi}{\mathbb{E}_{\pi,z\sim p(Z)}\left[\sum_{t=0}^{T-1}(\phi(s_{t+1})-\phi(s_{t}))^{T}z\right]}$
@@
We provide the proof in Appendix A. Therefore, the meaning of maximizing traveled language-distance is equivalent to maximizing the WDM between $s_{T}$ and $z$ under language-distance. Any off-the-shelf model-free RL algorithms can then be used to maximize the sum of rewards. We used PPO [Schulman et al., 2017] for as our primary RL algorithm because of its stable convergence. To ensure $\phi$ is 1-Lipschitz continuous under a distance metric $d$, we use dual gradient descent with a with a Lagrange multiplier $\lambda$ [Boyd and Vandenberghe, 2004].
4.2 Constraining the skill space via prompts
Figure 3: By prompting the LLM to generate different type of descriptions depending on the state, LGSD can adapt its focus during training.
The potential space for skills is vast, yet only a narrow region possesses semantic meanings. We want to explore only within a subspace which has inherent semantic meanings. To achieve this desiderata, we use language prompts to constrain the entire skill space into the desired semantic subspace.
@@ 4.2 Constraining the skill space via prompts
Moreover, we highlight that we can induce the LLM to not only focus on a single semantic aspect of the scene but also to adapt its focus onto different semantic aspects throughout the training. For instance, when the robot arm finally reaches the cube, the LLM is then asked to describe the object’s moved distance $d_{2}$. This additional prompt shifts the focus of the semantics to the object’s traveled distance and any other changes in the states are ignored by the LLM. This leads to the robot arm pushing the object to its maximum extent. Note that existing skill discovery algorithms have limited capability to constrain its focus by feeding subset of observations [Eysenbach et al., 2018, Park et al., 2023a]. In contrast, LGSD enables not only focusing on semantic dimensions but also adapting its focus across different semantics during training.

### v2

@@ Abstract
Skill discovery methods enable agents to learn diverse emergent behaviors without explicit rewards. To make learned skills useful for downstream tasks, obtaining a semantically diverse repertoire of skills is crucial. While some approaches use discriminators to acquire distinguishable skills and others focus on increasing state coverage, the direct pursuit of ‘semantic diversity’ in skills remains underexplored. We hypothesize that leveraging the semantic knowledge of large language models (LLM) can lead us to improve semantic diversity of resulting behaviors. In this sense, we introduce Language Guided Skill Discovery (LGSD), a skill discovery framework that aims to directly maximize the semantic diversity between skills. LGSD takes user prompts as input and outputs a set of semantically distinctive skills. The prompts serve as a means to constrain the search space into a semantically desired subspace, and the generated LLM outputs guide the agent to visit semantically diverse states within the subspace. We demonstrate that LGSD enables legged robots to visit different user-intended areas on a plane by simply changing the prompt. Furthermore, we show that language guidance aids in discovering more diverse skills compared to five existing skill discovery methods in robot-arm manipulation environments. Lastly, LGSD provides a simple way of utilizing learned skills via natural language.
@@ 1 Introduction
One of the key capabilities of intelligent agents is to autonomously learn useful skills applicable to downstream tasks without task-specific objectives. Consider how human infants acquire manipulation skills through simple play with toys, which can later be applied to a broader set of tasks, such as using a spoon or holding a bottle. Our research aims to emulate this process of skill learning. Prior works in unsupervised skill discovery suggest that maximizing diversity of behaviors can be a way to develop such skills (Gregor et al., 2016; Eysenbach et al., 2018; Kwon, 2020). These approaches operate under the assumption that acquiring a wide range of diverse behaviors may lead to the development of useful skills for downstream tasks. Building on this premise, studies like Sharma et al. (2019); Hansen et al. (2019); Liu & Abbeel (2021a); Laskin et al. (2022) have associated behaviors with random vectors by maximizing Mutual Information between them to acquire diverse skills. Additionally, there are other approaches aimed at increasing coverage of the state space (Liu & Abbeel, 2021b; Yarats et al., 2021; Park et al., 2021), or exploration-based strategies that leverage errors in predictive models to enhance learning (Burda et al., 2018; Pathak et al., 2019; Park et al., 2023a).
@@
However, we contend that these measures are proxies for what we term “semantic diversity” and they do not necessarily reflect the semantic diversity of a repertoire of skills. For instance, in high degrees of freedom (DOF) systems, such as robot-arm manipulation tasks, unstructured movements of each DOF might exhibit high diversity in terms of state coverage or distinctiveness from a neural discriminator, but nonetheless lack meaningful semantic diversity. To more accurately measure semantic diversity, we propose the use of Large Language Models (LLM) as they are well-suited for understanding and assessing the semantic significance of different behaviors in a way that better aligns with human judgment.
Figure 1: We proposed LGSD  which can discover a semantically distinctive set of skills. We showcase four sample skills acquired from a single training run. Our approach successfully learned skills that manipulate only ‘edible’ objects (banana and meat_can) from a total of four objects.
@@
In this work, we present Language Guided Skill Discovery (LGSD), a skill discovery algorithm that utilizes the guidance of LLMs to learn semantically diverse skills. LGSD aims to fulfill three desiderata for skill discovery. (i) Firstly, we want to discover skills that have diverse semantic meanings. We use LLMs to generate descriptions for each of agent states. Based on these descriptions, we measure the semantic difference between states using language-distance, and train skills to maximize this distance. Because LLMs are equipped with a semantic understanding of the world, this provides a way to maximizing semantic differences between skills. Fig. 1 illustrates skills learned from LGSD, showcasing the semantic diversity of resultant behaviors. (ii) Secondly, we seek to restrict the search space for skills into a desired semantic subspace. Intuitively, restricting the skill space is akin to enforcing an infant to grab toys only with their hands, and not with their mouth or feet. We implement this desideratum by utilizing language prompts. Humans provide prompts to LLMs, and the generated descriptions are then constrained by these prompts. This will allow our method to ignore differences outside the intended semantic space. (iii) Lastly, we aim for the learned skills to be easy to use. When the skill space is continuous, merely selecting the best skills for target tasks becomes non-trivial. LGSD supports a zero-shot language instruction-following capability by training separate network $\psi$ for skill inference. Humans can provide a natural language description of the desired state, and then LGSD can infer which skill should be used to reach that goal state.
@@ 2.1 Unsupervised skill discovery
Common approaches to associate skills with corresponding behaviors are based on maximizing the mutual information (MI) between states $S$ and skills $Z$, or $\mathcal{I}(S;Z)$ (Gregor et al., 2016; Eysenbach et al., 2018; Sharma et al., 2019; Hansen et al., 2019; Liu & Abbeel, 2021a; Laskin et al., 2022). MI can be viewed as a dependency measure that quantifies the association between two random variables. To maximize this quantity, a popular early formulation utilizes the decomposition $\mathcal{I}(S;Z)=\mathcal{H}(Z)-\mathcal{H}(Z|S)$, where $\mathcal{H}[\cdot]$ refers to Shannon entropy. Since directly computing $\mathcal{H}(Z|S)$ is intractable, Eysenbach et al. (2018) introduce a variational posterior $q_{\phi}(z|s)$ to optimize the lower bound. Similarly, CIC (Laskin et al., 2022) maximizes $\mathcal{I}((s,s^{\prime});Z)$, promoting distinct state transitions across different skills. However, a significant limitation of using MI as the objective arises from the use of Kullback–Leibler (KL) divergence to define MI:
@@
Maximization of MI using KL divergence often results in skills with less distinctive behaviors. This is because KL divergence is fully maximized when two densities have no overlap, posing a problem as there is no further incentive to distinguish the densities beyond this point. This is problematic because having no overlap between two densities does not necessarily mean two skills are noticeably distinctive. In practice, this becomes a severe shortcoming of DIAYN (Eysenbach et al., 2018), where its objective can be fully optimized as long as the discriminator $q_{\phi}$ can distinguish them perfectly. Neural networks can easily distinguish minor numerical differences, so the skills learned with this objective often have only these minute differences.
@@
To overcome these shortcomings, recent skill discovery works (He et al., 2022; Park et al., 2021; 2023a; 2023b) propose using the Wasserstein Dependency Measure (WDM,  Ozair et al. (2019)):
@@
An essential characteristic of $\mathcal{I_{W}}$ is that it must be defined within a metric space. Therefore, the choice of metric to measure the difference between skills comes to the gist of the algorithm since it ultimately governs the resulting behavior of each skills. LSD (Park et al., 2021) employs the Euclidean distance in state space as its metric. Maximizing $\mathcal{I_{W}}$ under this metric encourages agents to visit states that are are far apart in terms of Euclidean distances. Conversely, CSD (Park et al., 2023a) points out the problem of using Euclidean distance as the metric. They argue that maximizing Euclidean distance tends to focus on “easy” changes in the state space, such as moving the robot arm itself rather than the target object. Thus, CSD proposed to maximize “controllability-aware distance” instead. This metric measures the negative log likelihood of a transition dynamics $-\log{p(s^{\prime}|s)}$. Using this metric means the algorithm favors visiting rare transitions. LGSD adopts a similar approach but differentiates itself by employing a language-based distance metric within the WDM framework, to maximize the semantic difference.
@@ 2.2 Using language as guidance for behavior learning
Recent work has studied using semantic knowledge in large pre-trained models for control tasks in several ways. A popular approach is to use the common-sense reasoning capabilities of LLMs to produce high-level plans over known, existing skills through the interface of natural language (Sharma et al., 2021; Ahn et al., 2022; Singh et al., 2023; Huang et al., 2022; Ha et al., 2023) or generated programs (Liang et al., 2022; Zeng et al., 2022; Vemprala et al., 2024). In addition, LLMs have also been shown to be effective in directly producing low-level actions that can be executed by an agent (Wang et al., 2023; Kwon et al., 2023; Driess et al., 2023; Brohan et al., 2023). Our work does not use LLMs to generate actions at any level of abstraction but to provide a meaningful distance measure to guide learning towards discovering semantically meaningful skills.
Several works have used LLMs to guide agents to learn semantically meaningful behavior by using them to design reward functions for either a given desired task (Yu et al., 2023; Ma et al., 2023). This showcases LLMs’ capacity to understand specific physical embodiments assorting it with semantic knowledge to reason about given tasks. We note that these works aim to solve a single specific task, whereas our method aims to learn a repertoire of diverse behaviors without explicit supervision.
The work most related to ours is Du et al. (2023), which employs an LLM to guide exploration. Their approach uses the LLM to suggest plausible set of goals for the agent, maximizing the cosine similarity between the language embeddings of the goals and the current state. While they focus on using an LLM to reason with a low-dimensional, discrete set of high-level actions in complex environments like Crafter and Housekeep, our work focus on discovering meaningful behaviors in complex, continuous control spaces where actions (e.g. joint targets) are not as readily interpretable. Furthermore, their reward function, based on language similarity, lacks theoretical support as it is not a valid metric. In contrast, our method effectively handles the language-distance measure and provides a theoretical framework for its use. We elaborate on this in Section 4.3.
@@ 4.1 Algorithm
Our goal is for an agent to learn a corpus of skills that visits semantically different states. Discovering semantically distinct states requires 1) a way of measuring the semantic difference between states, and 2) a means of maximizing this difference measure. To evaluate the semantic differences between different behaviors, we leverage the power of LLMs. We will consider two states as substantially distinct if the language descriptions produced by LLMs are semantically different. As illustrated in Fig. 2, an LLM is queried to produce a natural language description for each state. This description links the state to its semantic meaning. We then measure the difference between these language descriptions. To measure the difference between two language descriptions, we leverage a pre-trained natural language embedding model, Sentence-Transformer (Reimers & Gurevych, 2019), to map the language descriptions into fixed-dimension real-valued vectors. This allows us to measure the difference between these vectors using cosine similarity. We refer to this difference measure as language-distance and denote it as $d_{\text{lang}}$.
@@
Spreading out in the latent space can be achieved by aligning the vector $\phi(s^{\prime})-\phi(s)$ with a latent skill vector $z$ sampled from a Gaussian distribution $z\sim\mathcal{N}(0,I)$. Because the vector $z$ is sampled isotropically, aligning $\phi(s^{\prime})-\phi(s)$ with $z$ fosters a diversity of behaviors depending on the sampled $z$. We use the reward $r(s,z,s^{\prime})=(\phi(s^{\prime})-\phi(s))^{T}z$ and maximize the cumulative reward to achieve this objective. Here, Park et al. (2023b) showed that maximizing the cumulative reward $r(s,z,s^{\prime})$ with a 1-Lipschitz continuous $\phi$ is equivalent to maximizing the WDM between the terminal state $S_{T}$ and $Z$:
$\operatorname*{arg\,max}_{\pi,\phi}\ {\mathcal{I_{W}}(S_{T};Z)}$ $\approx\operatorname*{arg\,max}_{\pi,\phi}{\mathbb{E}_{\pi,z\sim p(Z)}\left[\sum_{t=0}^{T-1}(\phi(s_{t+1})-\phi(s_{t}))^{T}z\right]}$
@@
We provide the proof in Appendix B. Therefore, the meaning of maximizing traveled language-distance is equivalent to maximizing the WDM between $s_{T}$ and $z$ under language-distance. Any off-the-shelf model-free RL algorithms can then be used to maximize the sum of rewards. We used PPO (Schulman et al., 2017) for as our primary RL algorithm because of its stable convergence. To ensure $\phi$ is 1-Lipschitz continuous under a distance metric $d$, we use dual gradient descent with a with a Lagrange multiplier $\lambda$ (Boyd & Vandenberghe, 2004).
4.2 Constraining the search space via prompts
Figure 3: By prompting the LLM to generate different descriptions depending on the state, LGSD can adapt its focus during training.
The potential space for skills is vast, yet only a narrow region possesses semantic meanings. We want to explore only within a subspace which has inherent semantic meanings. To achieve this desiderata, we use language prompts to constrain the entire search space into the desired semantic subspace.
@@ 4.2 Constraining the search space via prompts
Moreover, we highlight that we can induce the LLM to not only focus on a single semantic aspect of the scene but also to adapt its focus onto different semantic aspects throughout the training. For instance, when the robot arm finally reaches the cube, the LLM is then asked to describe the object’s moved distance $d_{2}$. This additional prompt shifts the focus of the semantics to the object’s traveled distance and any other changes in the states are ignored by the LLM. This leads to the robot arm pushing the object to its maximum extent. Note that existing skill discovery algorithms have limited capability to constrain its focus by feeding subset of observations (Eysenbach et al., 2018; Park et al., 2023a). In contrast, LGSD enables not only focusing on semantic dimensions but also adapting its focus across different semantics during training.